### v1

# Using shadow of black hole to examine the energy exchange between axion matter and rotating black hole

Abstract
We find that a slowly rotating axion-modified black hole resulting from the backreaction of an axion field on a rotating Kerr black hole can have a D-shaped shadow as that for a highly counter-rotating Kerr black hole. This attributes to the fact that the energy exchange between the axion matter and the black hole influences the rotation of the black hole, so the black hole angular momentum first decreases to zero and then the BH starts to rotate to the opposite direction. Further increasing the coupling leads to a “human-face-like” shaped shadows and new lensings due to the chaotic scattering, which are novel and drastically different from Kerr black hole. Our analysis provides the first counterexample to that slowly rotating black hole has nearly circular shadow.
I Introduction
In a realistic Universe the black holes (BHs) are rotating which are described by the Kerr metric Kerr:1963ud . For a rotating BH described by the Kerr metric, the description of the behaviour of matter is more complicated than in the non-rotating case. The metric in this case is only stationary, whilst the spherical symmetry has been replaced by axisymmetry. A further difficulty is the complexity of the coordinate systems for describing processes near a Kerr BH. The behaviour of matter in the neighbourhood of rotating BHs has been extensively studied. In the pioneering work, Bardeen, Press and Teukolsky had studied the physical processes around the BHs, together with their properties and interactions with the surrounding matter Bardeen:1972fi . To avoid the complexity in the perturbation analysis, in most cases it is assumed that the background metric is slowly rotating. Their method has essentially assumed that the non-rotating matter cancels out the “frame-dragging” effects of the BH rotation. In a sense, such a matter is co-rotating with the BH in such a way that the physical processes can be analyzed in their own frame. The authors of Bardeen:1972fi have used the method of locally non-rotating frames, which gives a clear picture of the process of energy extraction. The amount of such energy however, for realistic BHs, is not astrophysically significant.
In modified theories of gravity like the Chern-Simons (CS)-gravity theories with the action Jackiw:2003pm ; Zanelli:2005sa ; Alexander:2009tp ; Yunes:2009hc ; Yagi:2012ya
$\begin{split}S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa^{2}}-\frac{1}{2}(\partial_{\mu}b)(\partial^{\mu}b)-AbR_{CS}\right],\end{split}$ (1)
where $b$ is a pseudoscalar denoting the axion matter field, $A$ is the coupling parameter of the axion field to the gravitational CS term $R_{CS}$ with dimension of length, and $\kappa^{2}=8\pi G$ with $G$ is the Newton’s gravitational constant in four dimensional spacetime. Note that we shall use natural units $G=c=1$. The CS modification is a total derivative, and thus innocuous unless pseudoscalar (axion-like) fields couple to it, in which case it deforms non-trivially the gravitational field of spinning black holes. Such deformations are described by a modified Kerr geometry. These Kerr-like BH solutions contain second-order curvature corrections (fourth-order in derivatives), and also characterise the low-energy-limit of string theory Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz . In the context of string theory, the axion fields are part of the excitation spectrum, and they are of two types Svrcek:2006yi . One, is the so-called string-model independent axion, which in (3+1)-dimensions is dual Duncan:1992vz ; Svrcek:2006yi to the field strength of the so-called Kalb-Ramond, spin-one, antisymmetric-tensor field, which plays the role of a totally antisymmetric torsion in the geometry. The second type of axions in string theory arise from compactification of the extra dimensions. Potential signatures of CS theories in the emission of gravitational waves, produced in the inspiral of stellar compact objects into massive black holes, both for intermediate and extreme mass ratios, have been investigated in Sopuerta:2009iy . Also in Canizares:2012is , the authors attempted to answer the question as to what extent the extreme mass ratio inspirals observations with a space based gravitational wave observatory like LISA might be able to distinguish between CS theories and General Relativity (GR).
Rotating BHs have been studied extensively in such CS theories. First, they were found as solutions of second-order field equations for the metric stemming from string theories, in the presence of axions, in Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz , and more recently in Boulware:1985wk ; Gonzalez:2010vv . In particular, by extending non-trivially previous results Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz ; Yunes:2009hc ; Yagi:2012ya ; Myung:2020etf , slowly rotating BH solutions, in the presence of axion fields back reacting on the geometry, have been recently investigated Chatzifotis:2022mob . Its metric is
$\begin{split}ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{1}{\left(1-\frac{2M}{r}\right)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}{\theta}d\phi^{2}\right)-2r^{2}a\sin^{2}\theta W(r)dtd\phi~{},\end{split}$ (2)
with
$\begin{split}W(r)=\frac{2M}{r^{3}}-\frac{\gamma^{2}(189M^{2}+120Mr+70r^{2})}{14r^{8}}+O(A^{2n})~{},~{}~{}~{}~{}\gamma^{2}=\frac{A^{2}\kappa^{2}}{M^{4}}~{},\end{split}$ (3)
where $M$, $\gamma$ and $a$ are the mass of BH, coupling parameter, and rotation parameter, respectively. The back reaction of the axionic field on the slowly-rotating background metric has been studied to arbitrary order in perturbation theory of an appropriate parameter, $\gamma$, proportional to the axion-CS interaction coupling. The expansion generated by the perturbation theory allows one to go arbitrarily close to the horizon of the hairy (with axion-hair Campbell:1990ai ; Campbell:1990fu ; Duncan:1992vz ) BH. One can thus study the behaviour of the axionic matter outside the horizon of the slowly-rotating Kerr-type metric and see how this matter distribution influences the Kerr-like BH itself. In the case of the Kerr BH, the total angular momentum is given by $Ma$ Carter:1968rr . In this respect we stress that in Chatzifotis:2022mob , the presence of the axion field coupled to the CS term, which back reacts on the Kerr BH background, implies that the total angular momentum consists of the contributions of both the axion field and the hairy Kerr-like BH itself, which arises as a consequence of the axionic back reaction on the geometry.
In Chatzifotis:2022ene the study in Chatzifotis:2022mob of the axion-modified-Kerr BH system was extended by searching for potentially observable effects associated with bound trajectories of satellites in such backgrounds. As in Chatzifotis:2022mob a key role in the analysis will be played by the aforementioned perturbative parameter $\gamma$, which is proportional to the coupling of the axion field with the CS term and inversely proportional to the square of the total energy of the spacetime. This dimensionless parameter, rather than the coupling constant itself, that appears in the correction function of the modified Kerr-like BH solution and measures the strength of the back reaction on the geometry. The axion matter outside the horizon of the BH acquires an angular momentum in such a way that the total angular momentum of the system (BH and axionic matter) remains constant and equal to $Ma$. It was shown that as $\gamma$ increases, the BH angular momentum decreases and there exist a critical value $\gamma_{c}$ beyond which the BH angular momentum increases while the BH starts to rotate to the opposite direction. As such, the angular momentum of the BH horizon may in principle reach large values in magnitude, while the slowly rotating approximation remains valid with a total angular momentum equal to $Ma$. We understand this effect as a counterplay of two competing systems; the Kerr-like BH and the axionic matter rotating outside the horizon of the BH.
In this work we study possible observational effects, like shadow and lensing Perlick:2021aok ; Cunha:2018acu ; Wang:2022kvg ; Lukes-Gerakopoulos:2010ipp , of the axion-modified-Kerr BH with the metric (2). The coupling between the axion and gravity modifies the equations of motions in such a way as to allow an exchange of energy between the axion field and gravity. We expect to probe the evidence of energy exchange by BH shadow, and provide alternative BH shadow profile, since nowadays is golden era to study various phenomenological deviations from GR BH but has potential to be detected by Event Horizon Telescope. Throughout all physical quantities are rescaled by $M$ so that we will work with $M=1$.
II Shadow and Lensing by Axion induced Kerr-like black hole
In order to obtain the distorted apparent sky for a given observer, namely the shadow and lensing effect of Kerr-like BH, we shall compute photon geodesics in the above geometry (2). The equations of motion of the photons can be determined from the Hamilton-Jacobi equation,
$\mathcal{H}=-\frac{\partial S}{\partial\lambda}=\frac{1}{2}g_{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}}=0~{},$ (4)
where $\mathcal{H}$, $S$ and $\lambda$ are the canonical Hamiltonian, the Jacobi action and affine parameter respectively. For the geometry, the system does not have the symmetry associated with Carter constant and cannot be integrated over it. Therefore, we use the numerical backward ray-tracing method to investigate the lensing effect and shadow of BH. To this end, we employ the setup in Bohn:2014xxa ; Cunha:2015yba ; Hu:2020usx ; Zhong:2021mty , so we consider that the BH is illuminated by the ball-like source which is placed at far distance, and the BH and the camera are inside the ball-like source. The observer uses the fish-eye camera to photograph the BH.
As shown in Fig. 1(a), the ball-like extended source is divided into four symmetric parts, each part is marked with a different color, and the intersection of the four colors is marked with a white dot. The black longitudinal and latitudinal lines divide the ball-like source into the grids, and the interval between adjacent longitudinal or latitudinal lines is $\pi/18$. In order to investigate the trajectories of light rays, we can take the orthonormal tetrad at the observer’s position
$\begin{split}e_{0}=\frac{g_{\phi\phi}\partial_{t}-g_{\phi t}\partial_{\phi}}{\sqrt{g_{\phi\phi}(g^{2}_{\phi t}-g_{\phi\phi}g_{tt})}},~{}~{}~{}e_{1}=-\frac{\partial_{r}}{\sqrt{g_{rr}}},~{}~{}~{}e_{2}=\frac{\partial_{\theta}}{\sqrt{g_{\theta\theta}}},~{}~{}~{}e_{3}=-\frac{\partial_{\phi}}{\sqrt{g_{\phi\phi}}},\end{split}$ (5)
where $g_{\mu\nu}$ is the metric components of Kerr-like BH. Since the optical paths are reversible, the light rays reaching the observer can be considered to be emitted by the observer. These light rays have two endings, either returning to the light source or falling into the BH, where the light rays falling into the black hole correspond to the shadow of the BH.
(a)   Spherical source
(b)   Schwarzschild
(c)   Kerr, $a=0.999$
(d)   Kerr, $a=-0.999$
(e) $\gamma=0$, $a=0.2$
(f) $\gamma=1$, $a=0.2$
(g) $\gamma=1.5$, $a=0.2$
(h) $\gamma=2$, $a=0.2$
Figure 1: Shadow and Lensing by slowly rotating Kerr-like BHs with large enough coupling parameters. Here the inclination angle of the observer is fixed at $\theta_{o}=\pi/2$.
In Fig. 1, the images of Schwarzschild BH, Kerr BH $(a=0.999,-0.999)$, and Kerr-like BHs with different coupling parameters $\gamma$ are plotted. The observer is placed at $\theta_{o}=\pi/2$, the images have obvious symmetry, and the lensing effect can be clearly observed. The black area in the center is the shadow of the BH, and the white ring outside the shadow is the Einstein ring. The shadow of Schwarzschild BH is a perfect circle as can be seen in Fig. 1(b). Once we turn on the rotation in Kerr BH, $D$-shape will appear but it is faint, unless the spin $a$ is large enough, see for example $a=0.999$ (Fig. 1(c)) and $a=-0.999$ (Fig. 1(d)).
From the above figures it is easy to understand that for $a=0.2$ and $\gamma=0$, a tiny distortion exists in the left side. As we increase $\gamma$, the left distortion will become fainter because increasing $\gamma$ makes the BH angular momentum decrease but not yet change the rotating direction. There should be a critical $\gamma_{c}$ for which the angular momentum vanishes such that the shadow recover perfect circle. If we further increase $\gamma$ beyond $\gamma_{c}$, as shown in Chatzifotis:2022ene the BH angular momentum increases while it starts to rotate to the opposite direction, so the shadow distortion could appear in the other side because of the counter-rotating effect. This indeed explicitly happens in Fig. 1(h) which is similar to that in Kerr BH with $a=-0.999$ in Fig. 1(d). Therefore, starting with a slowly co-rotating BH with $a=0.2$ as we increase the coupling of the axion field to the CS term, because there is an exchange of energy between the axion field and the strong gravity, we observe a shadow which mimics that of a fast counter-rotating Kerr BH.
(a) $\gamma=2.5$, $a=0.2$
(b) $\gamma=3$, $a=0.2$
(c) $\gamma=4$, $a=0.2$
(d) $\gamma=5$, $a=0.2$
Figure 2: Shadow and Lensing by slowly rotating Kerr-like BHs with large enough coupling parameters. Here the inclination angle of the observer is fixed at $\theta_{o}=\pi/2$.
(a) $\gamma=2.5$, $a=0.2$
(b) $\gamma=3$, $a=0.2$
(c) $\gamma=4$, $a=0.2$
(d) $\gamma=5$, $a=0.2$
Figure 3: Shadow boundary of the slowly rotating Kerr-like BHs extracted from Fig. 2.
In Fig. 2 and Fig. 3, the shadows and lensing effect of the slowly rotating Kerr-like BH with larger coupling parameters $\gamma$ are plotted. Comparing to the Kerr BH, we observe two novel structures of the shadow for the Kerr-like BH with $a=0.2$ : (i) The deformation of shadows of the slowly rotating Kerr-like BH with positive $a=0.2$ is on the same side of Kerr BH with negative $a$. Moreover, though the whole spacetime rotates slowly, the distortion of the shadow is significant. This means that non-trivial interaction between the axion and gravity could construct a slowly rotating spacetime as a whole, but has an internal structure in which a highly rotating event horizon is allowed, leading to highly rotating effects. (ii) Instead of standard $D$-shaped shadow for Kerr BH, here we observe “human-face-like” shaped shadow. This may provide a new attempt for us to use very large baseline interferometry (VLBI) observations EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022apq of the shadows to verify the existence of axion matter. Moreover, two disconnected crescent-shaped shadows appeared directly above and below the “human-face-like” shaped shadow, which is more explicit in Fig.3. Besides, the lensing effects, in particular in the vicinity of the distorted shadow, are novel and different from Kerr BH. As we comparing the lensing from Fig. 2(a)-Fig. 2(d), we detect that a regime with pink and blue lights will merge on the symmetric axial near the distorted shadow curve, as the coupling $\gamma$ increases. It means that in this Kerr-like BH, some photons can have more then one radial turning point, generically corresponding to chaotic motion Zelenka:2017aqn ; Destounis:2021mqv ; Destounis:2020kss , which distinguish from only one radial turning point in Kerr BH Wilkins:1972rs . The disconnected shadows and lensings show that the interaction between the axionic matter and gravity, and the distribution of axion significantly affect the chaotic scattering, so does the chaotic region.
III Conclusion and discussion
Starting from an axion modified Kerr-like BH, we observe that a slowly rotating Kerr-like BH can produce a similar $D$-shaped shadow as that for a rapidly counter-rotating Kerr BH. These phenomena seems to be a natural consequence of the energy exchange between the axion field and BH such that a slowly rotating spacetime can also form a rapidly rotating event horizon. As the interplay becomes stronger, novel “human-face-like” shaped shadows and peculiar lensing phenomena, which significantly shift from those of Kerr BH, can be observed. Apparently, the rich lensing effects attributes to the chaotic scattering of the photons around the Kerr-like BH. An exhaustively studies on the bound orbit, light ring, and trapped trajectories Cunha:2016bjh could help us to obtain a deeper understanding of all those phenomenon, which shall be present elsewhere. In addition, it will be worthwhile to consider fast rotating Kerr-like BH to further study the effects among the fast rotation parameter $a$ and coupling parameter $\gamma$ on the shadows in the future.
Acknowledgements.
This work is partly supported by the Natural Science Foundation of China under Grants No. 12375054, Natural Science Foundation of Jiangsu Province under Grant No. BK20211601, and the Postgraduate Research &\&& Practice Innovation Program of Jiangsu Province under Grants No. KYCX22-3452.